OpenStudy (anonymous):

A boat on the potomac river travels the same distnace downstream in 2/3 hour as it does going upstream in 1 hr. If the speed of the curent is 3 mph, find the speed of the boat in the still water.

OpenStudy (phi):

you need to know a few things: 1) rate * time = distance 2) the speed going with the current is faster than going against the current if the current goes C and the boat goes B, going with the current, the speed is B+C going against the current, the speed is B-C

OpenStudy (phi):

current is 3 mph, so the speed with the current is B+3 the speed against the current is B-3

OpenStudy (phi):

downstream in 2/3 hour using rate * time = distance we write (B+3) * 2/3 = D we use the speed going with the current, and time = 2/3 hour

OpenStudy (phi):

going upstream in 1 hr write (B-3) * 1 = D using the speed going against the current, and time = 1 hour

OpenStudy (phi):

travels the same distance means we can say (B+3) * 2/3 = (B-3) * 1 can you solve for B ?

OpenStudy (anonymous):

No, still have no idea what I'm doing

OpenStudy (phi):

There are a few parts to this problem. One is finding the equations The second part is doing the algebra.

OpenStudy (phi):

We can go over solving the equation (without understanding how we got the equation) start with (B+3) * 2/3 = (B-3) * 1 and solve for B. first, notice on the right side you have (B-3)*1 you can simplify that to (B-3) (because multiplying by 1 does not change (B-3) ) you have \[ \frac{2}{3}(B+3) = B-3 \] can you multiply both sides by 3?

OpenStudy (anonymous):

Yeah and it will take out the 3 in the 2/3 and then multiply it by -3 on the other side so.. 2b+6=b-9

OpenStudy (phi):

almost. to multiply by both sides by 3, write 3 times each side, like this \[ 3\cdot \frac{2}{3}(B+3) = 3\cdot (B-3) \] on the left side 3/3 is 1 so that simplifies to \[ 2(B+3) = 3(B-3) \] can you simplify both sides... by distributing the 2 on the left, and 3 on the right side?

OpenStudy (anonymous):

2(B+3)=3(B-3) 2B+6=3b-9 then 2b-2b , and -2b from the 3b on the other side to get b 6=b-9 Then +9 and move it over to the 6 to get the b alone 15=B

OpenStudy (phi):

B stands for the Boat's speed. so the answer to find the speed of the boat in the still water. is 15 miles per hour